# From BERT to LLMs:Comparing and Understanding Chinese Classifier Prediction in Language Models

### v1

@@ Abstract
Ziqi Zhang, Jianfei Ma, Emmanuele Chersoni, Jieshun You, Zhaoxin Feng The Hong Kong Polytechnic University zhangziqi.bram@gmail.com, majianfei030706@gmail.com, emmanuelechersoni@gmail.com, jacksonyou666@outlook.com, BettyFeng2003@gmail.com
@@ 1 Introduction
In the NLP literature, there has been extensive exploration of Natural Language Understanding (NLU) using Pre-trained Language Models (PLMs) Wang et al. (2022) and Large Language Models (LLMs) Chang et al. (2024). However, all current studies on Chinese classifiers only include behaviours of traditional models for classifier prediction or selection with limited interpretability efforts  Peinelt et al. (2017); Järnfors et al. (2021), while to the best of our knowledge, there is no evaluation study with more recent autoregressive LLMs. In addition to the interest in evaluating LLMs on this essential component of Chinese grammar, it should be kept in mind that learning classifier systems has been proven to be particularly challenging for learners of Chinese as L2 Liang (2008); Liu (2018), and thus NLP technologies with a robust knowledge of classifiers would be a precious resource to develop educational tools.
@@ 2.4 Classifier Ranking by Log Probability
Log Probability (LogProb) is a popular quantitative metric to test models’ uncertainty by activating with given inputs Portillo Wightman et al. (2023); Kong et al. (2020). It is widely used in various tasks to reflect LMs’ uncertainty about inputs and generated content Jiang et al. (2021).
Tested by Hu and Levy (2023); Kauf et al. (2024), log probability also works well in LLMs, like detecting the grammar and the semantic plausibility of sentences, matching human behavior, and performing better than the outputs from directly prompting. However, it should be noted that this score is often sensitive to some spurious factors, including word frequency and output length Salazar et al. (2020); Holtzman et al. (2021).
Therefore, we insert classifier candidates into a sentence to compute the average LogProb of each filled sentence to avoid the influence of spurious factors. In theory, the best-fitting classifier should be the one with the highest LogProb sentence score.
@@ 3 Methodology
3.1 Dataset
We employ the Chinese Classifier Dataset Peinelt et al. (2017), a comprehensive resource with annotated classifier-noun pairs in sentential contexts, convenient to adapt classifier prediction tasks. This dataset contains 681,104 sentences, encompassing 172 distinct classifiers that nearly cover the entire commonly used Mandarin classifiers. Additionally, the Stanford constituent parser Levy and Manning (2003) was applied to annotate the head noun in each sentence. Due to computational resource limitations, we randomly sampled 11,917 sentences that span all classifiers and preserve their original distribution for further processing. These sampled instances were split into training and test sets at an 85:15 ratio to support fine-tuning and evaluation.
@@ 3.3 LLM-Based Classifier Prediction
With locally deployed Qwen3, we replace the empty classifier position (indicated with an underscore in the dataset) in each sentence with each of 172 candidate classifiers and use the IncrementalLMScorer from the minicons111minicons is a Python library for efficient probability scoring of transformer-based language models. Please refer to its’ github link: https://github.com/kanishkamisra/minicons to extract the log probability of each filled sentence by averaging the token-level log probabilities.It can be represented as:
@@
where $n$ is the total number of samples, and $rank_{i}$ is the rank of the correct classifier for the $i$-th sample within the model’s top 3 predictions (1, 2, or 3), or 4 if it is not in the top 3.
@@ 4.4 How LMs’ Attention Mechanisms Contribute to Prediction?
Given the classifiers’ strong dependence on their head nouns, the differences in attention mechanisms between BERT and LLMs, and the above analysis, we further investigate how the architectural distinctions account for the gaps. We select BERT as our baseline reference and employ 4 different attention masking types distinct from the standard attention masking for BERT MLM. Examples and comparisons are shown in  3.
@@ 6 Conclusions
Though this research provides insights on the comparative performance of BERT and large language models (LLMs) for Chinese classifier prediction, there are also limitations that should be acknowledged.
@@ A.1 Hardware
All experiments were conducted using a single NVIDIA A100 GPU with 40GB of dedicated memory and a single NVIDIA H20 GPU with 96GB of memory, hosted on a system equipped with an AMD EPYC 9K84 96-core processor (16 vCPUs) and 150 GB of system RAM. Each experimental run was configured to have a duration exceeding three hours.
@@ A.2 Experiement Setup
Transformers: version 4.55.2 (Hugging Face)
PyTorch: version 2.8.0
@@ A.3 Prompts Usage
We utilized the the prompt shown in table 6 to make our model inference and fine tune.
@@
Table 6: Prompts Introduction
@@ Appendix B Instruction and Statistics of Classifier Annotation
To assess the reliability and the objectivity of this annotation, we calculated the Inter-Annotator Agreement (IAA) through the application of Cohen’s kappa to a randomly selected sample of 500 classifier tokens. We thereby reached the IAA score of 82.68%, the existence of which indicates high annotator agreement and the robustness of our scheme of classifier categorization. For example, Event Classifier had been used in the counting of action or event occurrences. It does classify events and not physical objects. Therefore, the diction example like "场", "次", "趟" can be categorized into Event Classifier. Finally, we got the statistics about the the Count and Frequency for each classifier type, which can be seen in table 7. This table presents the frequencies of the six classifier types on the test set in numbers. The extreme imbalance of frequency often (e.g., Individual classifiers as the majority, Approximate/Standard classifiers as the minority) offers invaluable background in understanding the fine-grained model performance investigation by classifier type. It indicates the degree of difficulty and availability of data for each of the classifier types during the evaluation process.
Classifier Type Count Frequency

### v2

@@ Abstract
Ziqi Zhang, Jianfei Ma, Emmanuele Chersoni, Jieshun You, Zhaoxin Feng Department of Language Science and Technology, The Hong Kong Polytechnic University {blameredens.zhang, jianfei-mark.ma, jieshun.you, zhaoxinbetty.feng}@connect.polyu.hk, emmanuele.chersoni@polyu.hk
@@ 1 Introduction
In the natural language processing (NLP) literature, there has been extensive exploration of Natural Language Understanding (NLU) using Pre-trained Language Models (PLMs) Wang et al. (2022) and Large Language Models (LLMs) Chang et al. (2024); Ma et al. (2025). However, all current studies on Chinese classifiers only include behaviours of traditional models for classifier prediction or selection with limited interpretability efforts  Peinelt et al. (2017); Järnfors et al. (2021), while to the best of our knowledge, there is no evaluation study with more recent autoregressive LLMs. In addition to the interest in evaluating LLMs on this essential component of Chinese grammar, it should be kept in mind that learning classifier systems has been proven to be particularly challenging for learners of Chinese as L2 Liang (2008); Liu (2018), and thus NLP technologies with a robust knowledge of classifiers would be a precious resource to develop educational tools.
@@ 2.4 Classifier Ranking by Log Probability
Log Probability (LogProb) has been proven effective for various token-level tasks like assessing grammatical correctness and semantic plausibility, where its outputs often align with human judgments and outperform direct prompting Hu and Levy (2023); Kauf et al. (2024).
We acknowledge that this metric is not without limitations. It is known to be sensitive to confounding variables such as word frequency [^1] The analysis of effect for word frequency on accuracy is attached in Appendix D and output length Salazar et al. (2020); Holtzman et al. (2021). But they are difficult to isolate due to the uneven distribution nature of pre-training data. Despite this sensitivity, LogProb also be recognized as the robust choice for check the certainty of LMs’ output because its validity and superiority for semantic tasks are strongly supported by Kauf et al. (2024).
To specifically mitigate the influence of output length,we insert classifier candidates into a sentence to compute the average LogProb of each filled sentence to directly obtain the score without any redundant generation In this framework, the candidate that yields the highest sentence-level LogProb is considered the best fit.
@@ 3 Methodology
3.1 Dataset Constructions
We employ the Chinese Classifier Dataset Peinelt et al. (2017), a comprehensive resource with annotated classifier-noun pairs in sentential contexts, convenient to adapt classifier prediction tasks. This dataset contains 681,104 sentences, encompassing 172 distinct classifiers that nearly cover the entire commonly used Mandarin classifiers. Additionally, the Stanford constituent parser Levy and Manning (2003) was applied to annotate the head noun in each sentence. Although classifiers exhibit diverse pairings in pragmatic contexts, their syntactic component combinations are highly fixed, as supported by  Li and Thompson (1989), suggesting that a relatively small number of examples is sufficient to effectively evaluate their accurate usage and prediction. Due to this and computational resource limitations, we initially randomly sampled 11,986 sentences that span all classifiers and preserve their original distribution. After manual screening to remove 69 erroneous cases (e.g., annotation errors or syntactic anomalies), we obtained 11,917 valid sentences for further processing. These sampled instances were split into training and test sets at an 85:15 ratio to support fine-tune and evaluation.
@@ 3.3 LLM-Based Classifier Prediction
With locally deployed Qwen3, we replace the empty classifier position (indicated with an underscore) in each sentence with each of 172 candidate classifiers and use the IncrementalLMScorer from the minicons222minicons is a Python library for efficient probability scoring of transformer-based language models. Please refer to its’ github link: https://github.com/kanishkamisra/minicons to extract the log probability of each filled sentence by averaging the token-level log probabilities.It can be represented as:
@@
where $n$ is the total number of samples, and $rank_{i}$ is the rank of the correct classifier for the $i$-th sample within the model’s top 3 predictions (1, 2, or 3), or 4 if it is not in the top 3. The choice of a top-3 cutoff for R-rank was deliberate. Given that many Chinese nouns collocate with multiple classifiers, evaluating the top-3 candidates provides a sufficiently broad and practical assessment of a model’s discriminative ability. Extending this range further would dilute the metric’s practical significance, as lower-ranked candidates are increasingly unlikely to be contextually appropriate.
@@ 4.4 How LMs’ Attention Mechanisms Contribute to Prediction?
Given the classifiers’ strong dependence on their head nouns, the differences in attention mechanisms between BERT and LLMs, and the above analysis, we further investigate how the architectural distinctions account for the gaps. We select BERT as our baseline reference and employ 4 different attention masking types distinct from the standard attention masking for BERT MLM. Examples and comparisons are shown in Figure 3.
@@ 6 Conclusions
Though this research provides insights on the comparative performance of BERT and LLMs for Chinese classifier prediction, there are also limitations that should be acknowledged.
@@ A.1 Hardware
All experiments were conducted using a single NVIDIA A100 GPU with 40GB of dedicated memory and a single NVIDIA H20 GPU with 96GB of memory, hosted on a system equipped with an AMD EPYC 9K84 96‑core processor (16 vCPUs) and 150 GB of system RAM. Each experimental run was configured to have a duration exceeding three hours.
@@ A.2 Experiement Setup
This study involves Transformers packages with version 4.55.2 (Hugging Face) and platform of PyTorch with version 2.8.0.
@@ A.3 Prompts Usage
We utilized the the prompt shown in table 6 to make our model inference and fine tune. Due to the task is Chinese classifier prediction, we directly apply Chinese as the target language for prompt, the English version is also attached to the table for reference.
@@
Table 6: The demonstration of prompts used in both the inference and fine-tuning stages is provided below. The English version is a direct, literal translation of the primarily used Chinese version.
@@ Appendix B Instruction and Statistics of Classifier Annotation
To assess the reliability and the objectivity of this annotation, we calculated the Inter-Annotator Agreement (IAA) through the application of Cohen’s kappa to a randomly selected sample of 500 classifier tokens. We thereby reached the IAA score of 82.68%, the existence of which indicates high annotator agreement and the robustness of our scheme of classifier categorization. For example, Event Classifier had been used in the counting of action or event occurrences. It does classify events and not physical objects. Therefore, the diction example like "场", "次", "趟" can be categorized into Event Classifier. Finally, we got the statistics about the the Count and Frequency for each classifier type, which can be seen in Table 7. It presents the frequencies of the six classifier types on the test set in numbers. The extreme imbalance of frequency often (e.g., Individual classifiers as the majority, Approximate/Standard classifiers as the minority) offers invaluable background in understanding the fine-grained model performance investigation by classifier type. It indicates the degree of difficulty and availability of data for each of the classifier types during the evaluation process.
Classifier Type Count Frequency (%)
@@ Appendix C Chinese Classifier Categories and Explanation
Appendix D Relationship of classifier frequency and accuracy
Group NoC Count Proportion (%)
Group 1 1 677 34.95
Group 2 1 113 5.83
Group 3 1 102 5.27
Group 4 4 120 6.20
Group 5 10 116 5.99
Group 6 15 120 6.20
Group 7 15 120 6.20
Group 8 15 120 6.20
Group 9 15 120 6.20
Group 10 21 118 6.09
Group 11 40 120 6.20
Group 12 34 91 4.70
Table 8: This table demonstrates the rank based on the frequency of classifiers from the highest frequency of group 1 to lowest frequency group 12 and distribution of the classifiers. The NoC(number of categories) demonstrate how many categories of classifiers in this group.
Figure 4: Line chart of accuracy for used models in groups divided by classifier frequency from high to low. The dashed line represents the fine-tuned model, and the solid line represents the original model.
To explore the relationship between classifiers’ frequency and accuracy, we sorted and grouped the classifiers that appeared in Chinese Classifier Dataset Peinelt et al. (2017) based on their frequency of occurrence. With the grouping results summarized in Table 8, we first assign the classifier exhibiting the highest frequency to Group 1. The classifiers with the second and third highest frequencies are then placed into Group 2 and Group 3, respectively. To achieve a more balanced distribution across categories, and considering that Groups 2 and 3 are of appropriate scale to accommodate additional classifiers, we establish a count threshold of 120 for group assignment. Any remaining classifiers that meet this threshold are sequentially allocated to subsequent groups in the processing order. We then separately aggregated the results of all applied models on the test set according to these groupings, as shown in Figure  4. It can be observed that, overall, the accuracy of language models in predicting quantifiers decreases as the frequency of quantifiers in the dataset declines, indicating the prediction of classifier (or LogProb ranking) are influenced by the word frequency.